# Coxeter embeddings are injective

### v1

@@ Abstract.
We show that certain embeddings of Coxeter groups within other Coxeter groups are injective.
@@ 1. Coxeter partitions and Coxeter embeddings
    For each $s\in S$, the finite parabolic subgroup $\hat{W}_{U(s)}$ associated to $U(s)\subseteq\hat{S}$ is a finite Coxeter group. Let $w_{0}^{s}$ denote the longest element of $\hat{W}_{U(s)}$.
@@
Proposition 1.3.
The Coxeter embedding is a well-defined homomorphism.
Proof.
For $s\neq t\in S$, the element $\phi(s)\phi(t)=w_{0}^{s}w_{0}^{t}$ has order $m_{st}$ by definition. Any longest element is an involution, so $\phi(s)^{2}=\operatorname{id}$ for any $s\in S$. ∎
Our main theorem is as follows:
@@
Coxeter embeddings are injective, and send reduced expressions to reduced expressions. Moreover, for $w\in W$ and $\hat{s}\in\hat{S}$, letting $s=\pi(\hat{s})$, we have $\phi(w)\hat{s}<\phi(w)$ if and only if $ws<w$.
The rest of this section is devoted to proving this theorem.
We will use some standard notions and results for Coxeter groups, which can all be found (or deduced) from [BB05, §3 and §4]. From here on, the partial order $\leq$ (and $<$) on any Coxeter system refers to the weak right Bruhat order. We use the following notation. For elements $x,y,z$ in a Coxeter group, we write $x=y.z$ whenever $x=yz$ and $\ell(x)=\ell(y)+\ell(z)$; this is called a reduced composition. We write $\operatorname{RD}(w)$ for the right descent set of $w$, i.e. the set of simple reflections $s\in S$ such that $ws<w$. Recall that $\operatorname{RD}(w)$ always generates a finite Coxeter group. Moreover, if $x=y.z$ then $\operatorname{RD}(z)\subset\operatorname{RD}(x)$.
If $(W,S)$ is a Coxeter system and $J\subseteq S$, write $W^{J}$ for the set of minimal length representatives for cosets in $W/W_{J}$. If $y\in W^{J}$ and $z\in W_{J}$ then $y.z$ is reduced. Any $x\in W$ has a unique decomposition $x=y.z$ with $y\in W^{J}$ and $z\in W_{J}$.
@@
Remark 1.6.
The proof of Theorem 1.4 is modelled on [Cri99, Lemma 2.2]. However, we note that Lemma 2.2, as stated in loc. cit., is actually false. Namely, an LCM-homomorphism $\phi:A^{+}_{\Gamma}\rightarrow A^{+}_{\hat{\Gamma}}$ does not guarantee that $\phi$ sends square-free elements to square-free elements (which allows one to deduce that the induced map on the corresponding Coxeter groups sends reduced expressions to reduced expressions). Consider the Artin monoids associated to the following Coxeter graphs:
{tikzpicture}
The homomorphism defined by $\phi(s)=s_{1}s_{2}$ and $\phi(t)=t_{1}t_{2}$ is an LCM-homomorphism, but it does not send square-free elements to square-free elements, since
$\phi(sts)=s_{1}s_{2}t_{1}t_{2}s_{1}s_{2}=s_{2}t_{1}s_{2}t_{2}s_{1}s_{1}$
is not square-free. The issue is that when $st\in A^{+}_{\Gamma}$ has no common multiple, an LCM-homomorphism imposes no condition on the images of $\langle s,t\rangle^{m}$. Condition 2(b) in our definition of Coxeter partitions was added to handle exactly this situation.
2. Examples of Coxeter embeddings
@@
We present here some examples of Coxeter partitions of interest, which include (and generalize) some examples that were considered by others.
@@
sends reduced expressions to reduced expressions (e.g. via Lemma 2.2 below). We prove that this remains true even when $\hat{W}$ is infinite, by showing that Lusztig’s partitions are a specific type of Coxeter partitions. For this we will require two results about reduced expressions for powers of Coxeter elements. The first lemma is classical, and the second is a result of Speyer.
Lemma 2.2 ([Bou02, Chapter V §6 Exercise 2]).
@@
Proposition 2.4.
@@
By Lemma 2.3, any power of $\hat{c}$ is reduced [^1] In fact, we only need a weaker version of Lemma 2.3, which was proven earlier in [FZ07, Corollary 9.6].. But for all $k\geq 2$, $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is always a subword of some power of $\hat{c}$, so $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ must itself be reduced.
@@
If each $\hat{W}_{i},$ has the same Coxeter number $m_{st}$ (finite or infinite), then by the connected case above, $\langle w_{0}^{s,i},w_{0}^{t,i}\rangle^{k}$ is reduced for $2\leq m\leq m_{st}$. As a (direct) product of reduced expressions, $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is also reduced, which is condition (2) of a Coxeter partition.
@@
Throughout, we will implicitly use the fact that a Lusztig’s partition is a Coxeter partition, shown in Proposition 2.4.
@@
Coxeter embeddings associated to Lusztig’s partitions (when each $m_{st}$ is finite) were studied by Lusztig in [Lus83, §3]. These include examples coming from folding by graph symmetries, such as the first example below. Lusztig’s famous inclusion of Coxeter groups of type $H_{4}$ into type $E_{8}$ is a Lusztig’s partition that does not come from graph symmetries. Our setting of Lusztig’s partitions includes examples where $m_{st}$ is infinite (see Example 2.8).
@@
{tikzpicture}
The reader is encouraged to construct the affine-type examples given in [Cri99, Table 2].
@@
{tikzpicture}

### v2

# mastercopy.bib

@@ Abstract.
We show that certain embeddings of Coxeter groups within other Coxeter groups are injective using the notion of Coxeter partitions. Moreover, we study Lusztig’s partitions, which are generalizations of Lusztig’s admissible maps and Crisp’s foldings. We show that they classify the simplest type of Coxeter partitions, whose embeddings of Coxeter groups send each generator to a product of commuting generators. Consequently, these embeddings are also injective, and we prove that they preserve Coxeter numbers. These results were previously known, due to work of Mühlherr and Dyer.
Authors’ notes
This short paper contains no novel results. When we first wrote this paper we believed Theorems 1.3 and 2.4 were new, after having asked experts and failing to find it in the literature. Subsequently, thanks to some helpful leads, we eventually tracked down the original sources. We will leave this paper on the arXiv to make the results more accessible, but do not intend to publish it. Discussion of prior work can be found in §4.
@@ 1. Coxeter partitions and Coxeter embeddings
    For each $s\in S$, the parabolic subgroup $\hat{W}_{U(s)}$ associated to $U(s)\subseteq\hat{S}$ is a finite Coxeter group. Let $w_{0}^{s}$ denote the longest element of $\hat{W}_{U(s)}$.
@@
It is straightforward to verify that the Coxeter embedding is well-defined.
Our first result is as follows:
@@
Coxeter embeddings are injective, and send reduced expressions to reduced expressions. Moreover, for all $w\in W$ and $\hat{s}\in\hat{S}$, setting $s=\pi(\hat{s})$, we have $\phi(w)\hat{s}<\phi(w)$ if and only if $ws<w$.
This theorem is originally due to Mühlherr [Muhlherr_CoxinCox], see §4; we provide a short proof of the theorem below. Examples of Coxeter partitions can be found in §3, and a special family of Coxeter partitions satisfying $\hat{W}_{U(s)}\cong\prod_{i}\mathbb{Z}/2\mathbb{Z}$ (i.e. a product of type $A_{1}$ finite Coxeter groups) for all $s\in S$ will be studied in §2.
We will use some standard notions and results for Coxeter groups, which can all be found (or deduced) from [Comb_Coxeter_book, §2, §3 and §4]. From here on, the partial order $\leq$ (and $<$) on any Coxeter system refers to the weak right Bruhat order. We use the following notation. For elements $x,y,z$ in a Coxeter group, we write $x=y.z$ whenever $x=yz$ and $\ell(x)=\ell(y)+\ell(z)$; this is called a reduced composition. We write $\operatorname{RD}(w)$ for the right descent set of $w$, i.e. the set of simple reflections $s\in S$ such that $ws<w$. Recall that $\operatorname{RD}(w)$ always generates a finite Coxeter group. Moreover, if $x=y.z$ then $\operatorname{RD}(z)\subset\operatorname{RD}(x)$.
If $(W,S)$ is a Coxeter system and $J\subseteq S$, write $W^{J}$ for the set of minimal length representatives for cosets in $W/W_{J}$. Note that $W^{J}$ is the set of $w\in W$ such that $\operatorname{RD}(w)\cap J=\emptyset$. If $y\in W^{J}$ and $z\in W_{J}$ then $y.z$ is reduced. Any $x\in W$ has a unique decomposition $x=y.z$ with $y\in W^{J}$ and $z\in W_{J}$.
@@
2. Lusztig’s partitions
@@
sends reduced expressions to reduced expressions (e.g. via Lemma 2.2 below). We prove that this remains true even when $\hat{W}$ is infinite, by showing that Lusztig’s partitions are a specific type of Coxeter partitions. In fact, we will also show that Lusztig’s partitions are precisely Coxeter partitions which satisfy (the stronger) condition (1) of Definition 2.1. In other words, Lusztig’s partitions classify the “simplest” type of Coxeter partitions, where the elements of $U(s)$ are all pair-wise disjoint in the Coxeter graph of $(\hat{W},\hat{S})$. These will all be the content of Theorem 2.4.
We begin by recalling two lemmas about reduced expressions for powers of Coxeter elements. The first lemma is classical, and the second is a result of Speyer.
Lemma 2.2 ([Bourbaki_Lie_4-6, Chapter V §6 Exercise 2]).
@@
Our second result is the following theorem, originally due to Dyer [Dyer_rootII], see §4.
Theorem 2.4.
@@
By Lemma 2.3, any power of $\hat{c}$ is reduced [^1] In fact, we only need a weaker version of Lemma 2.3, which was proven earlier in [FZ_clustercoeff, Corollary 9.6].. But for all $k\geq 2$, $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is always a subword of some power of $\hat{c}$, so $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ must itself be reduced.
@@
If each $\hat{W}_{i}$ has the same Coxeter number $m_{st}$ (finite or infinite), then by the connected case above, $\langle w_{0}^{s,i},w_{0}^{t,i}\rangle^{k}$ is reduced for $2\leq m\leq m_{st}$. As a (direct) product of reduced expressions, $\langle w_{0}^{s},w_{0}^{t}\rangle^{k}$ is also reduced, which is condition (2) of a Coxeter partition.
@@
Throughout, we will implicitly use the fact that a Lusztig’s partition is a Coxeter partition, shown in Theorem 2.4.
@@
3. Examples
Coxeter embeddings associated to Lusztig’s partitions (when each $m_{st}$ is finite) were studied by Lusztig in [LusSqint, §3]. These include examples coming from folding by graph symmetries, such as the first example below. Lusztig’s famous inclusion of Coxeter groups of type $H_{4}$ into type $E_{8}$ is a Lusztig’s partition that does not come from graph symmetries. Our setting of Lusztig’s partitions includes examples where $m_{st}$ is infinite (see Example 3.3).
@@ 3. Examples
[…]
The reader is encouraged to construct the affine-type examples given in [CrispInjective, Table 2].
@@
[…]
Remark 3.8.
Given a Coxeter system $(\hat{W},\hat{S})$ with $s_{+},s_{-}\in\hat{S}$ such that $\hat{m}_{s_{+}s_{-}}=3$, an edge contraction [li_heckecontract, §1.3] can be performed (on the Coxeter graph) to obtain a new Coxeter system $(W,S)$ defined by
$S$ $:=\left(\hat{S}\cup\{s_{0}\}\right)\setminus\{s_{+},s_{-}\};$
$m_{st}$ :=$:=\begin{cases}\hat{m}_{st},&\text{ if }s,t\neq s_{0};\\ \hat{m}_{ss_{+}}+\hat{m}_{ss_{-}}-2,&\text{ if }t=s_{0},\hat{m}_{ss_{+}}=2\text{ or }\hat{m}_{ss_{-}}=2;\\ \infty,&\text{ if }t=s_{0},\hat{m}_{ss_{+}},\hat{m}_{ss_{-}}>2.\end{cases}$
The group homomorphism $\phi:W\rightarrow\hat{W}$ sending $s_{0}\mapsto s_{+}s_{-}s_{+}$ with the identity on all other generators was shown to be injective in [li_heckecontract, Proposition 1.4.1]. While $\phi$ feels similar to a “Coxeter embedding”, we note that the natural surjection $\pi:\hat{S}\rightarrow S$ defined by $s_{+},s_{-}\mapsto s_{0}$ and identity on all other generators is not a Coxeter partition (condition (2) need not be satisfied). In fact, even though $\phi$ is injective, it need not send reduced expressions to reduced expressions (while a Coxeter embedding always does); this can already be seen from an edge contraction on the $A_{3}$ Coxeter graph.
4. Related works
We discovered after this paper was originally posted that none of our results (nor perhaps the proofs) are novel. We chose to keep our exposition and proofs intact above, and discuss this prior work here.
Admissible partitions; after Mühlherr
Our definition of Coxeter partitions turns out to be equivalent to Mühlherr’s definition of admissible partitions; see [Muhlherr_CoxinCox, pg. 278] for the definition and [Muhlherr_CoxinCox, Theorem 1.2 and Lemma 3.3] for the fact that they are equivalent (see also [Castella_admissiblesubmonoid, Lemma 14]).
The injective part of our Theorem 1.3 is [Muhlherr_CoxinCox, Theorem 1.1], whereas the preservation of reduced expressions is [Muhlherr_CoxinCox, Proposition 4.5 A1]. The final statement of Theorem 1.3 is in fact equivalent to Condition (A) of admissible partitions. Our proof of Theorem 1.3 above is different from (and shorter than) Mühlherr’s proof, and we now give context on that proof.
LCM-homomorphisms and Artin (sub)monoids
Oblivious to the earlier work by Mühlherr, our original intention was to use the results of LCM-homomorphisms (between Artin monoids) from Crisp’s work [CrispInjective] to obtain a weaker version of Theorem 1.3; cf. [CrispInjective, Proposition 2.3]. In the process, we also realized that the definition of LCM-homomorphisms and the proof of [CrispInjective, Lemma 2.2] (or similarly, [godelle_injectif, Proposition 2.6]) could be modified to obtain a direct proof of Theorem 1.3 in the Coxeter group setting, skipping the theory of Artin monoids.
The modification [^3] Condition 2(a) of Coxeter partitions is essentially the definition of LCM-homomorphisms in the free-of-infinity setting considered by Crisp in [CrispInjective], and Condition 2(b) was what we added to include the free-of-infinity setting. On the other hand, Godelle’s condition (L3) in his version of LCM-homomorphisms [godelle_injectif, Definition 2.1] for the general setting is strictly stronger than condition 2(b). required led us to the definition of Coxeter partitions, which we later found out is equivalent to Mühlherr’s definition of admissible partitions. We later found out that the generalization of both the works [CrispInjective, godelle_injectif] to the setting of admissible partitions (for Artin monoids) was studied by Castella in [Castella_admissiblesubmonoid]. Indeed, in Remark 35 of loc. cit. Castella suggests that the proofs of Crisp and Godelle for LCM-homomorphisms would also work in the more general setting of admissible partitions. Unwittingly, our proof of Theorem 1.3 is the realization of this remark, having stripped away the theory of Artin monoids.
Dyer’s work
Meanwhile, Dyer in [Dyer_rootII, Section 8] proves a generalization of Thoerem 1.3 (citing Mühlherr), and also proves Thoerem 2.4 in this generalized context. His proof of Theorem 2.4 is extremely similar to ours, including the use of Speyer’s result (Lemma 2.3). Dyer too seems to have been unaware of prior work of Crisp or Castella or Godelle.
Acknowledgements
The authors would like to thank Federica Gavazzi for pointing out the work of Godelle, which led us down the rabbit hole to the work of Castella then Mühlherr and finally Dyer. We also thank Yiqiang Li for informing us of his interesting work. B.E. was supported by NSF grant DMS-2201387, and his research group was supported by DMS-2039316.
[…]